OpenStudy (anonymous):

If A person puts one penny in a piggy bank on the first day, two pennies in on the second day, Four pennies in on the third day, eight pennies in on the fourth day,and so forth, how much money will be in the bank after 60 Days?

OpenStudy (ikram002p):

u can form a series such that S=1+2+3+8+.... now u need to find a formula of (a_n) so it could help u to expect the next 60 turm of the series.

OpenStudy (anonymous):

Would this be correct? a sub n = n *2^n?

OpenStudy (ikram002p):

mmm no it cant be , its true for the first turm but it dosnt goes well wid the second and th third

OpenStudy (anonymous):

I'm at a loss for words. My brain is exploding.

OpenStudy (rsadhvika):

\(a_n = n!!\)

OpenStudy (ikram002p):

will lets try this formula assuming (a_0=0) a_n= (a_n-1 * a_n-2) + a_2

OpenStudy (ikram002p):

now try it for the first 3 turm & forth turm u can see that 3= 2*1 +1 8= 3*2 +2

OpenStudy (anonymous):

is this progression correct 1, 2, 3???, 8 and not 1, 2, 4, 8

OpenStudy (anonymous):

I'm sorry. You are correct.

OpenStudy (anonymous):

It would be 1,2,4,8

OpenStudy (anonymous):

LOL, OK will have to rework then ... back in a few

OpenStudy (rsadhvika):

then its simple sequence of squares

OpenStudy (rsadhvika):

\(a_n = 2^{n-1}\)

OpenStudy (ikram002p):

@zJicez which formla is incorect !

OpenStudy (rsadhvika):

n = 1, 2, 3, ....

OpenStudy (anonymous):

Ikram I wrote out the problem wrong. I'm sorry, I edited the question.

OpenStudy (ikram002p):